In an economy, two goods: A and B are purchased in given fixed quantities for the prices shown below. Year Price of A $ Quantity of A Price of B $ Quantity of B 2016 16.00 25 2.00 10 2017 19.40 25 4.00 10 2018 22.80 25 6.00 10 Compute the CPI for each of the 3 years. Show your work.


1
Expert's answer
2021-03-12T09:16:21-0500

Total price of basket for each year is as shown below

"2016=16\times 25+2\times 10=420"

"2017=19.40\times 25+4\times 10=525"

"2018=22.80\times 25+6\times 10=630"

2016 is the base year

"CPI=" price of basket in year x "\/" price of basket in base year "\times 100"

"CPI_{2016}=420\/420=100"

"CPI_{2017}=525/420\times 100=125"

"CPI_{2018}=630/420\times 100=150"

inflation rate=(PI for year X)- (PI for base year)"\/" (PI for base year) "\times 100"

inflation rate year 2017"=(125-100)/(100)\times 100=25\%"

inflation rate year 2018"=" "=(150-100)/(100)\times 100=50\%"